Trace the conicoid represented by
x^2 +2z^2 = y. Also describe its sections by the planes x = c,∀c ∈ R.
1
Expert's answer
2020-05-18T18:50:06-0400

"x^2+2z^2=y"

Let's rewrite equation: "2x^2+4z^2=2y\iff \frac{x^2}{0.5}+\frac{z^2}{0.25}=2y"

Now we can see that our surface is elliptic paraboloid.


Denominators are unequal. It means that our paraboloid is not a paraboloid of revolution.


Let "x=0", then "y=2z^2"

And let "z=0", then "y=x^2"

"y=2z^2" and "y=x^2" principal parabolas.



Let's look on sections.

"y=c", where "c\isin R"

"x^2+2y^2=c -" ellipses if "c>0"



If "x=c," where "c\in R"

"y=2z^2+c^2 -" parabolas.

"c^2 -" is distance between the origin and vertex of parabola.



A similar result will be if we look at "z=c," where "c\in R"

"y=x^2+2c^2"